mr.stone started from home on a trip, planning to average 48 miles per hour. How fast must his son carl plan to travel in order to overtake him in 3 hours if carl started 30 minutes after his father?
Answer by mananth(16949)

mr.stone started from home on a trip, planning to average 48 miles per hour. How fast must his son carl plan to travel in order to overtake him in 3 hours if carl started 30 minutes after his father?
stone speed = 48mph
in 3 hours he travels 48*3=144 miles
..
let son's speed be be x mph more than Stne's speed
..
Son's speed = 48+x
..
Father starts 1/2 hour early
catch up distance = 24 miles
catch up time = 3 hours
time = distance / speed
3=24+144/(48+x)
3x+144 = 168
3x= 168-144
3x=24
x=8
...
son has to travel 48+8 =56 mph.